Tech Staff Summary 1/8/01
Current Priorities in addition to routine tasks
Eastern hard rock site has been identified and Mac has sent paperwork to owners to get agreement started.
Need to schedule a maintenance run.
Still need permissions on Edith and Library towers.
No word yet from Falcon.
Lawyers caused us to miss the January window. Mitch will call Mataxas to see where the agreement is.
Brewer needs to get the mockup done.
Need to have at the very least MKTA, KOSE, and Richardson ready to align when Library is installed. Greg is working on subcontract for KOSE, LNXT, and NMAD.
Greg and Jim will get the final supplies needed.
Brewer will install new wave_server disk so we can get MTTN stations on helicorder images.
John is waiting on callback from cable company for Cumberland Gap.
John will contact property owners at PKNC to try to reduce the annual fee from $350 to $300. Otherwise, we will go ahead and make the agreement. Hopefully this won't happen at too many other places.
Steve found out that LRNC reconfigured and cut off our feed. He is working with the local IS person to get it back.
John will call Steve to coordinate Athens alternatives. Need line of site with Star Mountain and some kind of continuous connection.
Getting reimbursed for the stolen truck and equipment is likely a wash. Mitch will check into getting vehicles insured either through an off campus rider or through private insurance.
Stays where it is. Need to patch holes in ceiling, add ventilation, yank non-functional shower, and install eye wash station.
This is becoming a high priority. We don't want to drag our feet but neither do we want to "burn" good field weather.
Will keep pluggin away as time permits.
Mitch still needs to complete the code to handle the gain channel. Subroutines are done and offline PII style is available. On the fly will need less robust than offline to avoid potentially large and time intensive buffering. (shall we start a pool on whether this ever gets done?)
Waiting on Schweig to clear out the rest of his junk.
Need to schedule training day with Bodin.
Pinnacle Towers now owns the NBC Bank roof rights and are actively buying up towers nationwide. Mitch and Greg will work on the agreement to NBC which hopefully won't require lawyers. Greg and Jim will visit NBC to assess what is there and plan for repeating PWLA.
On hold until we get USGS input. Greg will visit Pink Palace for their advice so that we can have a starting point when the GS people get here.
This is becoming a time sync. Mitch will work with Greg and Sharam to make sure Greg does not have his schedule exploded by it.
Second tier priorities
Mitch will get a plan together to permit the 20 new sites. These need
to be permitted by May 4, and operational by September 30.
Jim will install a map mosaic covering areas where CERI has hardware
Mitch will produce a station map for the ANSS-MA region
Chris will install the white board
Greg has found a supplier to fix the nasty old map.
Cannabalize PI to make as much as PII operational.
Check with Arch or Jer-Ming for account numbers to buy supplies.
EBZ/PWLA telemetry. Another repeater?
Secure propane tank.
replace coax jumpers
replace rg-58 unshielded jumpers
Need to record the SOH channel from the gps clocks to know when sync is lost (they will continue to produce IRIGE based on internal clock).
Greg has added channel to monitor gps clock. Need an ew module that monitors and complains.
Mitch and Steve will install earthworm v5 in June. This will include the
full blown schema capable of creating seed.
Status channels. Currently supplied on last 8 channels.
LNXT node permit/install L7 and L8
MKTA node permit/install remaining 3 ISIS